OpenStudy (anonymous):

factor by grouping x^3+4x^2-8x-32

OpenStudy (nenadmatematika):

x^2(x+4)-8(x+4)=(x+4)(x^2-8)=(x+4)(x-sqrt(8))(x+sqrt(8))